# Static Three-Dimensional Structures Determine Fast Dynamics Between Distal Loci Pairs in Interphase Chromosomes

### v2

@@ II Results
We then calculated the relative change in diffusivity as a function of the relative change in centrality. Fig. 7(b) shows that the centrality of chromatin loci decreases after cohesin deletion. Loci exhibiting a greater reduction in centrality typically show a larger percentage increase in diffusivity. Next, we investigate the two-point loci dynamics by calculating the relaxation time $\tau$. Fig. 7(c) shows a heatmap of the relative change in relaxation time between $\Delta$RAD21 and WT cells, $\ln(\tau_{\Delta\mathrm{RAD21}}/\tau_{\mathrm{WT}})$. Fig. 7(c) shows that although chromatin loci dynamics are accelerated after cohesin deletion, the change in two-point relaxation time $\tau$ is not uniform but is locus-dependent. For loci located within the TADs, the relaxation time increases after cohesin deletion because the distances between the loci increase. In contrast, loci located at the boundaries of TADs, the relaxation time decreases after cohesin deletion. These predictions are amenable to experimental tests.
We also computed the mode spectrum—characterized by the eigenvalues $\lambda_{p}$—for both wild-type (WT) and cohesin-depleted cells. Supplementary Figure 9(b) shows that the cohesin-depleted system exhibits a distinct gap between the $p=1$ and $p=2$ modes, suggesting a separation of relaxation time scales between the entire region and a subset of the region. This gap likely arises from the disruption of two loop extrusion domains present in the WT. Upon cohesin depletion, these domains disappear, and the separation between the $p=1$ and $p=2$ modes vanishes. Moreover, we observed a qualitative change in the scaling behavior of $|\lambda_{p}|$ with respect to $p$: in WT cells, $|\lambda_{p}|\propto p^{1.2}$, whereas in the cohesin-depleted cells, $|\lambda_{p}|\propto p^{1.5}$. The steeper scaling observed in the cohesin-depleted system indicates that small-scale fluctuations dissipate more rapidly, suggesting a loss of coordinated motion across larger chromatin domains.
@@ III Conclusion
Our theory, containing no adjustable parameter, accurately reproduces the experimental observations of loci dynamics in Drosophila embryo cells, thus resolving the apparent discordance between static chromatin structures and dynamic behaviors highlighted in recent studies [36]. Strikingly, the two-point relaxation times between chromatin loci scale with genomic separation is in excellent agreement with experiments demonstrating that the unexpected rapid relaxation dynamics are a consequence of effective long-range interactions, which could be mediated by factors like transcription factors and cohesin. By analyzing the eigenvalues of the connectivity matrix, we uncovered that the rapid chromatin dynamics exhibit complex, length-scale-dependent behavior, reflecting the hierarchical structural organization of chromosomes. This finding suggests that the dynamics of chromosomes cannot be captured by homopolymer models but requires knowledge of the intricate network of interactions in chromatin. We also used our model to predict the mean first passage times for contact between chromatin loci and found quantitative agreement with results calculated using the experimentally measured loci trajectories [36]. This calculation further supports the notion that chromatin dynamics are intrinsically linked to the precise static three-dimensional structure, which likely play a crucial role in processes such as promoter-enhancer communication.
@@
Additionally, our theory predicts that the heterogeneous single-locus diffusion behavior is dependent on local chromatin density. We found that loci with higher contact probabilities with other loci tend to exhibit slower diffusion, highlighting the influence of the interaction landscape on chromatin mobility. Our exploration of the effects of cohesin deletion revealed that chromatin loci in cohesin-depleted cells exhibit higher diffusivity and the changes in the two-point relaxation times are locus-dependent. This observation underscores the role of cohesin in constraining chromatin dynamics and organizing chromatin structure.
In summary, this work reinforces the concept that the static three-dimensional organization of chromosomes dictates their dynamic behavior, which resolves the conundrum raised in the experiments [36]. Importantly, we have shown that measurements of the contact map or the distance map using the imaging method are sufficient to calculate the loci-dependent chromatin dynamics. By establishing a direct and quantitative link between chromatin structure and dynamics, our general computational framework opens new avenues for exploring chromatin dynamics in various biological contexts.
Code Availability
The code for the model presented in this work and its detailed user instructions can be accessed at the GitHub repository: https://github.com/anyuzx/HIPPS-DIMES. The data analysis is performed using Python 3.12 in Jupyter Lab. The Python packages used in analyzing data are Scipy, Numpy, and Pandas.
@@ Acknowledgement
We are grateful to Mauro L. Mugnai for useful comments and discussions. This work was supported by a grant from the National Science Foundation (CHE 2320256) and the Welch Foundation through the Collie-Welch Chair (F-0019).

### v3

@@ II Results
Next, we investigate the two-point loci dynamics by calculating the relaxation time $\tau$. Fig. 6(c) shows a heatmap of the relative change in relaxation time between $\Delta$RAD21 and WT cells, $\ln(\tau_{\Delta\mathrm{RAD21}}/\tau_{\mathrm{WT}})$. Fig. 6(c) shows that although chromatin single-locus dynamics are accelerated after cohesin deletion, the change in two-point relaxation time $\tau$ is not uniform but is locus-dependent. For loci located within the TADs, the relaxation time increases after cohesin deletion because the distances between the loci increase. In contrast, loci located at the boundaries of TADs, the relaxation time decreases after cohesin deletion. These predictions are amenable to experimental tests. Mechanistically, cohesin depletion disrupts loop-extrusion–mediated insulation at TAD borders, allowing boundary-spanning loci to come into proximity and relax more quickly, while loci within TADs lose compaction and relax more slowly. Finally, we computed the mean first-passage time to contact, $\langle\tau_{c}\rangle$, as a function of genomic separation $s$ (Fig. 6(d)). In WT cells, loci within the same TAD obey $\langle\tau_{c}\rangle\sim s^{0.5}$, but pairs crossing TAD boundaries exhibit a sharp jump to much larger $\tau_{c}$. In contrast, in cohesin-depleted $\Delta$RAD21 cells $\langle\tau_{c}\rangle$ scales as $s^{1.2}$, with virtually no distinction between within-TAD and outside-TAD contacts — consistent with the loss of loop-extrusion–mediated insulation. This suggests that loop extrusion greatly reduces the time required for loci to contact within TADs.
We also computed the mode spectrum—characterized by the eigenvalues $\lambda_{p}$—for both wild-type (WT) and cohesin-depleted cells. fig. S10(b) shows that the cohesin-depleted system exhibits a distinct gap between the $p=1$ and $p=2$ modes, suggesting a separation of relaxation time scales between the entire region and a subset of the region. This gap likely arises from the disruption of two loop extrusion domains present in the WT. Upon cohesin depletion, these domains disappear, and the separation between the $p=1$ and $p=2$ modes vanishes. Moreover, we observed a qualitative change in the scaling behavior of $|\lambda_{p}|$ with respect to $p$: in WT cells, $|\lambda_{p}|\propto p^{1.2}$, whereas in the cohesin-depleted cells, $|\lambda_{p}|\propto p^{1.5}$. The steeper scaling observed in the cohesin-depleted system indicates that small-scale fluctuations dissipate more rapidly, suggesting a loss of coordinated motion across larger chromatin domains.
@@ III Discussion
Our theory, with no locus-specific fitting parameter and containing a single adjustable parameter — the effective friction coefficient that sets the overall time scale — accurately reproduces the experimental observations of loci dynamics in Drosophila embryo cells, thus resolving the apparent discordance between static chromatin structures and dynamic behaviors highlighted in recent studies [36]. Strikingly, the two-point relaxation times between chromatin loci scale with genomic separation is in excellent agreement with experiments demonstrating that the unexpected rapid relaxation dynamics maybe a consequence of effective long-range interactions, which could be mediated by factors like transcription factors and cohesin. By analyzing the eigenvalues of the connectivity matrix, we uncovered that the rapid chromatin dynamics exhibit complex, length-scale-dependent behavior, which may be connected to the hierarchical structural organization of chromosomes. This finding suggests that the dynamics of chromosomes cannot be captured by homopolymer models but requires knowledge of the intricate network of interactions in chromatin. A concise analytical connection linking eigenvalue spectrum to the dynamic scaling exponents remains an important avenue for future theoretical work. We also used our model to predict the mean first-passage times for contact between chromatin loci and found quantitative agreement with results calculated using the experimentally measured loci trajectories [36]. This calculation further supports the notion that chromatin dynamics are intrinsically linked to the precise static three-dimensional structure, which likely play a crucial role in processes such as promoter-enhancer communication.
@@
Additionally, our theory predicts that the heterogeneous single-locus diffusion behavior is dependent on local chromatin density. We found that loci with higher contact probabilities with other loci tend to exhibit slower diffusion, highlighting the influence of the interaction landscape on chromatin mobility. Our exploration of the effects of cohesin deletion revealed that chromatin loci in cohesin-depleted cells exhibit higher diffusivity and the changes in the two-point relaxation times are locus-dependent. This observation underscores the role of cohesin in regulating chromatin dynamics.
Although our main analyses focus on Drosophila chromosome 2R, we have also applied HIPPS-DIMES to human HCT116 cells to study cohesin depletion, demonstrating its direct applicability to mammalian chromosomes. In principle, the framework is agnostic to species or cell line. To underscore the generality of HIPPS-DIMES, we have also applied the theory to human GM12878 Hi-C data (Supplementary Materials and fig. S14) and to mESC Micro-C data at the Fbn2 locus (fig. S15), with comparable success in reproducing both single-locus and two-point dynamics. However, its current implementation treats each chromosome as a separate, contiguous polymer. Modeling on the whole-genome level would require either independent per-chromosome analyses or an extension to capture inter-chromosomal contacts.
It is important to recognize that HIPPS-DIMES relies on experimental data (Hi-C/Micro-C or imaging) as input, so any noise or error in these data will propagate into our predictions. Sources of error include uneven mappability and fragment-level biases in Hi-C/Micro-C, fixation artifacts introduced by Hi-C/Micro-C and many chromosome-tracing protocols (e.g., formamide-based FISH) [57], and localization uncertainty in imaging data [58, 59, 60]. To assess the impact of such errors, we verified that our dynamical predictions remain robust to moderate Hi-C/Micro-C noise and to realistic imaging localization errors ($\leq$ 50 nm); see Supplementary Materials and figs. S16–17. In future work, it would be valuable to extend our framework to explicitly model localization uncertainty or to apply HIPPS-DIMES to data generated by non-denaturing methods such as RASER-FISH [61, 62, 63].
In summary, by moving beyond homopolymer models like the Rouse and fractal globule models, which predict relaxation exponents that are inconsistent with experiments and fail to capture the rapid locus-dependent dynamics, we show that the precise, heterogeneous 3D structure inferred by HIPPS-DIMES dictates the observed chromatin dynamic behavior. The proposed theoretical framework, which could be applied to other systems, resolves the conundrum raised in the experiments [36]. Importantly, we have shown that measurements of the contact map using Hi-C/Micro-C or the distance map using the imaging method are sufficient to calculate the loci-specific chromatin dynamics. Because HIPPS-DIMES predicts time-resolved trajectories for chromatin loci, the predictions can be directly validated using live-cell tracking methods—such as CRISPR/dCas9 tagging, operator-repeat arrays, or MS2/MCP reporters—by comparing properties such as mean-square displacements, relaxation times, and first-passage statistics. By establishing a direct and quantitative link between chromatin structure and dynamics, our general theoretical framework opens avenues for exploring chromatin dynamics in various biological contexts.
@@ Acknowledgement
We are grateful to Mauro L. Mugnai for useful comments and discussions.
Funding: This work was supported by a grant from the National Science Foundation (CHE 2320256, recipient D.T.) and the Welch Foundation through the Collie-Welch Chair (F-0019, recipient D.T.).
Author contributions: G.S. and D.T. designed the research. G.S. performed the research. G.S., S.S. and D.T. analyzed the data. G.S., S.S., and D.T. wrote the manuscript.
Competing Interests: The authors declare no competing interests.
Data and Materials Availability: We used the following publicly available data: Micro-C data for Drosophila embryos are available from ArrayExpress (E-MTAB-9784; https://www.ebi.ac.uk/biostudies/arrayexpress/studies/E-MTAB-9784), multiplexed FISH imaging data for HCT116 cell line from GitHub (https://github.com/BogdanBintu/ChromatinImaging), Hi-C data for human GM12878 cells from ENCODE (ENCFF065LSP; https://www.encodeproject.org/files/ENCFF065LSP/), and Micro-C data for mESCs from GEO (GSE286495). The live-cell imaging data from Bruckner et al. [36] is publicly available at Zenodo repository [64] at https://zenodo.org/records/7616203. To process Hi-C/Micro-C data in .cool and .hic formats, we used the Cooler Python package [65] at https://github.com/open2c/cooler and the hic-straw Python package [66] at https://github.com/igvteam/hic-straw, respectively. The code for the model presented in this work and its detailed user instructions can be accessed at the Zenodo repository [67] at https://zenodo.org/records/15611946 or at GitHub repository for the latest version: https://github.com/anyuzx/HIPPS-DIMES. The data analysis is performed using Python 3.12 in Jupyter Lab. The Python packages used in analyzing data and visualization are Scipy, Numpy, Pandas, and Matplotlib. All data needed to evaluate the conclusions in the paper are present in the paper and the Supplementary Materials.
Supplementary Materials
  • Supplementary Text
  • Figs. S1 to S17